Plot a rating curve of flow rate (Q,cfs) vs. water surface elevation (ft) for an ogee spillway that is 35 ft long with a design height of 3.5 ft as the water surface elevation changes from 54 ft to 58 ft (at least 2 data points per foot). You may assume the weir crest height is 53.5 ft. **Your water surface elevation is equivalent to the sum of your h (that changes) and your weir crest height (constant).**
Answer #1

The water surface elevation = (58-54)+53.5 = 4+53.5 i.e 57.5ft.

Crest Length = 35ft

Crest Elevation = 57.5ft.

Therefore, Q= Cross Sectional Area* sqrt(2gH1) =35 * 57.5* sqrt(2*9.81*3.5)

= 35* 57.5*sqrt 68.87

= 16,701 cu.ft/sec.

Flow at H2 = 35* 57.5* sqrt(2*9.81*4)

= 17,828.5 cu.ft/sec

Therefore, curve